OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form. Find the exact value of the radical expression in simplest form. @Mathematics

OpenStudy (asnaseer):

where is the radical expression?

OpenStudy (anonymous):

One second im trying to insert the square root symbol

OpenStudy (asnaseer):

sounds painful :0)

OpenStudy (asnaseer):

so this is the question?\[\sqrt{p^6}\]

OpenStudy (anonymous):

\[\sqrt{p^6} + \sqrt{9p^6} + 3\sqrt{p^6} - \sqrt{p^6}\]

OpenStudy (asnaseer):

ok, use the following facts:\[\sqrt{a}=a^{1/2}\]\[(a^{n})^{m}=a^{nm}\]you then be able to find the answer

OpenStudy (agreene):

Or you could just see what cancels in that mess.. look for like terms and factor the 9p^6 one.

myininaya (myininaya):

\[\sqrt{p^6}+\sqrt{9}\sqrt{p^6}+3 \sqrt{p^6}-\sqrt{p^6}=\sqrt{p^6}(1+\sqrt{9}+3-1)\]

OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form.\[\sqrt{4y^3} - \sqrt{y^3}\]

OpenStudy (asnaseer):

yes, agreene is right - first simplify the expression - as myininaya has just done :-)

myininaya (myininaya):

\[\sqrt{p^6}=p^\frac{6}{2}=p^3\]

OpenStudy (agreene):

\[\sqrt{y^3}\]

OpenStudy (asnaseer):

use:\[\sqrt{ab}=\sqrt{a}*\sqrt{b}\]

OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form.\[2\sqrt{7x} +\sqrt{7x}- 3\sqrt{7x?}-7\]

OpenStudy (agreene):

-7

OpenStudy (asnaseer):

you should have all the information you need now to be able to solve these types of problems.

myininaya (myininaya):

\[(2+1-3)\sqrt{7x}-7\]